OpenStudy (anonymous):

When there are 30 girls in a dance, there are 12 boys in the dance. When there are 45 girls in the dance, there are 17 boys in the dance. Assume the relationship is linear. Which equation models the relationship between the number of girls, x, and the number of boys, y? y = 3x + 26 y=1/2x+22 y=1/3x+2 y=3x+9 Can someone please explain this? I don't want an answer.

OpenStudy (anonymous):

@ganeshie8 @johnweldon1993 @mathstudent55 @Destinymasha @undeadknight26

OpenStudy (nfcfox):

Let's start by finding the slope. We know that girls are x and that boys are y. We have two points already because they say when our x is 30 or 45 our y will be 12 or 17. The points would just be (30,12),(45,17)

OpenStudy (nfcfox):

To find slope we subtract our 1st y point by the 2nd y point all divided by the subtraction of our 1st x point by our 2nd y points. Therefore 17-12/45-30 17-12=5 and 45-30= 15. 5/15= 1/3

OpenStudy (anonymous):

OK, I knew that. What next? :)

OpenStudy (nfcfox):

At this point we can already tell what the answer is but we should find our y intercept nonetheless. We plug in our x slope and y into our point slope form. We can use either of the 2 points. 12=1/3(30)+B Where b is the y intercept.

OpenStudy (nfcfox):

12=10+B B=2 Therefore Our equation for the line in point slope form will be y= 1/3x +2

OpenStudy (nfcfox):

@CrazyCountryGirl do you understand it? It's just the equation of a line.

OpenStudy (anonymous):

Yeah, it makes sense. :) I learned some of it this morning in class, but the teacher went to fast. :)Thanks!!